2/5x + 10=0 what does x equal?
Answer by checkley77(12844)

2/5x + 10=0
2x/5=-10
2x=-10*5
2x=-50
x=-50/2
x=-25 ans.
Proof:
2*-25/5+10=0
-50/5+10=0
10-10=0
0=0